# Principled Federated Random Forests for Heterogeneous Data

### v1

@@ A.3 Candidate thresholds with quantile sketching
The proof proceeds in two steps: first establishing the error bound for local approximations, and then extending it to the global aggregate.
@@
Step 1: Local Error Bound. Consider a single client $k$ transmitting $B$ exact quantiles $q_{0},\dots,q_{B}$ corresponding to ranks $0,1/B,\dots,1$. For any value $x$ falling in the interval $q_{i},q_{i+1}]$, the monotonicity of the CDF implies that the true rank $F_{\nu,k}(x)$ lies within the rank interval $[i/B,(i+1)/B]$. Similarly, the linear interpolant $\hat{F}_{\nu,k}(x)$ constructed by the server is also bounded strictly within $[i/B,(i+1)/B]$. Since both the true function and the approximation are confined to the same interval of height $1/B$, their absolute difference is bounded by that height:
@@
Thus, the global rank error is strictly bounded by $\epsilon=1/B$, independent of the heterogeneity of the local distributions $F_{\nu,k}$. ∎
@@
We consider a fixed node $A$ and a centralized candidate $(j,t_{\mathrm{cent}})\in\mathcal{T}_{\mathrm{cent}}$. We omit the feature index $j$ in the notation and write $F_{\nu}$ and $\tilde{F}_{\nu}$ for the pooled empirical CDF and its reconstruction (as in Theorem 3.1). By construction, the federated candidate set for this feature is
@@
A.4 Communication reduction methods
A.4.1 Top-$L$ maximum gains aggregation recovers the true optimal feature
Proposition A.1 (Screening consistency under i.i.d. clients).
@@ A.4.1 Top-$L$ maximum gains aggregation recovers the true optimal feature
Appendix B Impurity Decomposition under Covariate Shift
@@
Appendix C Exact Midpoint Enumeration
For completeness, we describe how centralized CART split candidates can be reproduced exactly in a federated setting. At a given node $A$, centralized CART evaluates midpoints between consecutive unique feature values. In principle, this can be achieved federatively by having each client transmit its node-level feature values, which the server merges into a globally sorted list.
Formally, for feature $j$, let $\{v_{j,(r)}\}_{r=1}^{m_{A,j}}$ denote the sorted unique pooled values at node $A$. The exact candidate set is
$\mathcal{T}_{\nu}^{\mathrm{all}}=\bigcup_{j=1}^{d}\left\{(j,t):t=\tfrac{v_{j,(r)}+v_{j,(r+1)}}{2},\;1\leq r<m_{A,j}\right\}.$
@@ Appendix C Exact Midpoint Enumeration
Appendix D Additional Simulations
@@
D.1 Covariate shift: when local objectives mis-rank global splits
@@ D.1.1 Disjoint supports
D.1.2 Overlapping supports: failure driven by the heterogeneity term
Disjoint supports are not required for AvgImp to fail. Here, both clients have overlapping feature support, but the pooled impurity differs from the weighted average of local impurities due to the heterogeneity term $\mathcal{E}(A)$ in Theorem 3.3. We simulate $K=2$ clients with
@@ D.1.2 Overlapping supports: failure driven by the heterogeneity term
and a shared conditional model that linear in feature $X^{(1)}$:
@@
The pooled optimal split is near the pooled population mean (equal to $\mathbb{E}[X]=\frac{1}{2}(-\gamma)+\frac{1}{2}\gamma=0$ under balanced mixture weights), and every client has data around that point. Nevertheless, as $\gamma$ grows, $\mathcal{E}(A)$ increases, so maximizing the average local gain can favor thresholds near the client means (around $\pm\gamma$) instead of the pooled optimum. Table 3 confirms this effect: FedForest Quantiles remains close to centralized performance across shifts, while AvgImp degrades as $\gamma$ increases, despite overlapping supports.
@@
D.2 Covariate and Outcome Shift
We consider the fully heterogeneous regime of Equation 3, where both covariates ($\alpha_{k}=.5$ and $\gamma_{k}=3$) and outcomes ($\delta=1.5$) vary across clients. Unless stated otherwise, all methods are trained with a common random-forest configuration: $T=50$ trees, maximum depth $8$, minimum leaf size $5$, stratified (by client) bootstrap, and feature subsampling $\sqrt{d}$. We do not center $Y$ locally. When a method supports splitting on the site indicator, we enable Fisher-style optimal grouping of $H$.
Data are generated with $K=5$ clients and $n_{k}=200$ samples per client in dimension $d=10$.

### v2

@@ C.1 Candidate thresholds with quantile sketching
The proof proceeds in first establishing the error bound for local approximations, and then extending it to the global aggregate.
@@
Step 1: Local Error Bound. Consider a client $k$ transmitting $B$ exact quantiles $q_{0},\dots,q_{B}$ corresponding to ranks $0,1/B,\dots,1$ of a given feature (we omit the feature index $j$ in the notation) at node $\nu$. $\forall x\in[q_{i},q_{i+1}]$, the monotonicity of the CDF implies that the true rank $F_{\nu,k}(x)\in[i/B,(i+1)/B]$. Similarly, the linear interpolant $\hat{F}_{\nu,k}(x)$ constructed by the server is also bounded strictly within $[i/B,(i+1)/B]$. Since both the true function and the approximation are in the same interval of height $1/B$, their absolute difference is bounded by that height:
@@
Thus, the global rank error is strictly bounded by $\epsilon=1/B$, independent of the heterogeneity of the local distributions $F_{\nu,k}$.
@@
We consider a centralized candidate $(j,t_{\mathrm{cent}})\in\mathcal{T}_{\mathrm{cent}}$. We write $F_{\nu}$ and $\tilde{F}_{\nu}$ for the pooled empirical CDF and its reconstruction (as in Theorem 3.1). By construction, the federated candidate set for this feature is
@@
C.2 AvgImp impurity reduction estimator: homogeneous setting
Proof of Theorem 3.3.
By Equation 5, the approximation error equals the split heterogeneity correction exactly:
$\left|\Delta I(j,t;\mathcal{S}_{\nu})-\widehat{\Delta I}_{\mathrm{AvgImp}}(j,t;\mathcal{S}_{\nu})\right|=|\Delta\mathcal{E}(j,t;\nu)|,$
with $\Delta\mathcal{E}(j,t;\nu):=\mathcal{E}(\nu)-\frac{n_{L}}{n_{\nu}}\mathcal{E}(\nu_{L})-\frac{n_{R}}{n_{\nu}}\mathcal{E}(\nu_{R}),$ so it suffices to show $\mathcal{E}(\nu)=O_{p}(K/n_{\nu})$ for any node $\nu$.
Regression (variance impurity).
Let $\mu_{\nu,k}:=\frac{1}{n_{\nu,k}}\sum_{i\in\mathcal{S}_{\nu,k}}Y_{i}$ and $\mu_{\nu}:=\frac{1}{n_{\nu}}\sum_{i\in\mathcal{S}_{\nu}}Y_{i}$ denote the local and global empirical means at node $\nu$, respectively. The node heterogeneity is $\mathcal{E}(\nu)=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}(\mu_{\nu,k}-\mu_{\nu})^{2}\geq 0$. Under 2.1, each $\mu_{\nu,k}$ is an unbiased estimator of the population mean $\mu:=\mathbb{E}[Y]$, and so is $\mu_{\nu}$, giving $\mathbb{E}[\mu_{\nu,k}-\mu_{\nu}]=0$ and $\mathrm{Var}(\mu_{\nu,k})=O(1/n_{\nu,k})$ (finite by $\mathbb{E}[Y^{2}]<\infty$). Taking expectations:
$\mathbb{E}\bigl[\mathcal{E}(\nu)\bigr]=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}\,\mathbb{E}\bigl[(\mu_{\nu,k}-\mu_{\nu})^{2}\bigr]\leq\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}\,\mathrm{Var}(\mu_{\nu,k})=\sum_{k=1}^{K}O\!\left(\frac{1}{n_{\nu}}\right)=O\!\left(\frac{K}{n_{\nu}}\right),$
where we used $\mathbb{E}[(\mu_{\nu,k}-\mu_{\nu})^{2}]\leq 2\,\mathrm{Var}(\mu_{\nu,k})+2\,\mathrm{Var}(\mu_{\nu})=O(1/n_{\nu,k})$. Since $\mathcal{E}(\nu)\geq 0$, Markov’s inequality gives $\mathcal{E}(\nu)=O_{p}(K/n_{\nu})$.
Classification (Gini and entropy impurities).
Let $\mathbf{p}_{\nu,k}:=\bigl(p_{\nu,k}^{(c)}\bigr)_{c=1}^{C}$ and $\mathbf{p}_{\nu}:=\bigl(p_{\nu}^{(c)}\bigr)_{c=1}^{C}$ be the local and global empirical class-probability vectors at node $\nu$, where $p_{\nu,k}^{(c)}:=\frac{1}{n_{\nu,k}}\sum_{i\in\mathcal{S}_{\nu,k}}\mathbf{1}[Y_{i}=c]$ and $p_{\nu}^{(c)}:=\frac{1}{n_{\nu}}\sum_{i\in\mathcal{S}_{\nu}}\mathbf{1}[Y_{i}=c]$ for each class $c\in\{1,\ldots,C\}$.
For the Gini index, $\mathcal{E}_{\mathrm{Gini}}(\nu)=\sum_{k}\frac{n_{\nu,k}}{n_{\nu}}\|\mathbf{p}_{\nu,k}-\mathbf{p}_{\nu}\|^{2}$, and since $\mathbb{E}\|\mathbf{p}_{\nu,k}-\mathbf{p}_{\nu}\|^{2}=O(1/n_{\nu,k})$ by the same moment argument, $\mathcal{E}_{\mathrm{Gini}}(\nu)=O_{p}(K/n_{\nu})$.
For entropy, $\mathcal{E}_{\mathrm{Entropy}}(\nu)$ equals the Jensen–Shannon divergence
$\mathrm{JSD}\bigl(\{\mathbf{p}_{\nu,k}\}\,\|\,\mathbf{p}_{\nu}\bigr):=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}D_{\mathrm{KL}}\bigl(\mathbf{p}_{\nu,k}\,\|\,\mathbf{p}_{\nu}\bigr),\qquad D_{\mathrm{KL}}(\mathbf{p}\,\|\,\mathbf{q}):=\sum_{c=1}^{C}p^{(c)}\log\frac{p^{(c)}}{q^{(c)}}.$
For each $k$, a second-order Taylor expansion of $D_{\mathrm{KL}}(\mathbf{p}_{\nu,k}\|\mathbf{p}_{\nu})$ around $\mathbf{p}_{\nu,k}=\mathbf{p}_{\nu}$ gives
$D_{\mathrm{KL}}\bigl(\mathbf{p}_{\nu,k}\,\|\,\mathbf{p}_{\nu}\bigr)=\frac{1}{2}\sum_{c=1}^{C}\frac{(p_{\nu,k}^{(c)}-p_{\nu}^{(c)})^{2}}{p_{\nu}^{(c)}}+O\bigl(\|\mathbf{p}_{\nu,k}-\mathbf{p}_{\nu}\|^{3}\bigr)=O\bigl(\|\mathbf{p}_{\nu,k}-\mathbf{p}_{\nu}\|^{2}\bigr),$
where the last equality holds provided $p_{\nu}^{(c)}\geq\epsilon>0$ for all $c$ (so the denominator is bounded away from zero). Hence $\mathcal{E}_{\mathrm{Entropy}}(\nu)=O_{p}(K/n_{\nu})$ by the same moment argument as for Gini.
Conclusion.
The same bound applies to the child nodes $\nu_{L}$ and $\nu_{R}$ (which each receive a $\Theta(n_{\nu})$ fraction of samples under a balanced-split regularity condition), so $|\Delta\mathcal{E}(j,t;\nu)|\leq\mathcal{E}(\nu)+\frac{n_{L}}{n_{\nu}}\mathcal{E}(\nu_{L})+\frac{n_{R}}{n_{\nu}}\mathcal{E}(\nu_{R})=O_{p}(K/n_{\nu})$.
C.3 Impurity Decomposition
Proof of Theorem 3.4 which states:
$\Delta I(j,t;\mathcal{S}_{\nu})=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}\Delta I(j,t;\mathcal{S}_{\nu,k})+\Delta\mathcal{E}(j,t;\nu),$
where the split-dependent heterogeneity term $\Delta\mathcal{E}(j,t;\nu)$ represents the change in the baseline node heterogeneity $\mathcal{E}(\nu)$ induced by the split $(j,t)$:
$\Delta\mathcal{E}(j,t;\nu)=\mathcal{E}(\nu)-\left(\frac{n_{L}(j,t)}{n_{\nu}}\mathcal{E}(\nu_{L}(j,t))+\frac{n_{R}(j,t)}{n_{\nu}}\mathcal{E}(\nu_{R}(j,t))\right).$
The baseline node heterogeneity $\mathcal{E}(\nu)$ term is defined as the Jensen gap of the local empirical target distributions:
where $\hat{P}_{k}$ is the empirical distribution of client $k$’s outcome at the node, and $\Psi$ is a standard statistical impurity function (e.g., variance, Gini index, or entropy–see Table 4).
For these split criteria, $\mathcal{E}(\nu)\geq 0$, with $\mathcal{E}(\nu)=0$ if and only if the local empirical outcome means are equal for all clients.
Proof.
We first establish the decomposition of the empirical impurity $I(\mathcal{S}_{\nu})$ at a single node $\nu$. Let $\mathcal{S}_{\nu}$ be the set of samples at node $\nu$, and $\mathcal{S}_{\nu,k}$ the subset of samples belonging to client $k$, such that $\mathcal{S}_{\nu}=\bigcup_{k=1}^{K}\mathcal{S}_{\nu,k}$ and $\mathcal{S}_{\nu,k}\cap\mathcal{S}_{\nu,k^{\prime}}=\emptyset$ for $k\neq k^{\prime}$. Let $P_{\nu}$ denote the empirical distribution of the target variable $Y$ over the pooled sample $\mathcal{S}_{\nu}$, and $P_{\nu,k}$ denote the empirical distribution over the local sample $\mathcal{S}_{\nu,k}$. The empirical distribution assigns mass $1/n$ to each observation. Since the datasets are disjoint, the global empirical measure $P_{\nu}$ is the weighted average of the local measures. For any event $E$:
$P_{\nu}(E)$ $=\frac{1}{n_{\nu}}\sum_{i\in\mathcal{S}_{\nu}}\mathbb{1}(y_{i}\in E)$
$=\sum_{k=1}^{K}w_{\nu,k}P_{\nu,k}(E),$
The global empirical impurity is defined as a function $\Psi$ (mean squared error, Gini, cross-entropy, etc.) applied to the global distribution $P_{\nu}$:
$I(\mathcal{S}_{\nu})$ $=\underbrace{\sum_{k=1}^{K}w_{\nu,k}\Psi(P_{\nu,k})}_{\text{Weighted Local Impurities}}+\underbrace{\left[\Psi\left(\sum_{k=1}^{K}w_{\nu,k}P_{\nu,k}\right)-\sum_{k=1}^{K}w_{\nu,k}\Psi(P_{\nu,k})\right]}_{\text{Jensen-gap / Heterogeneity Term }\mathcal{E}(\nu)}$
$=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}I(\mathcal{S}_{\nu,k})+\mathcal{E}(\nu).$
Next, we apply this single-node decomposition to the impurity reduction $\Delta I(j,t;\mathcal{S}_{\nu})$ of a candidate split $(j,t)$, which produces left and right children $\nu_{L}$ and $\nu_{R}$ of sizes $n_{L}$ and $n_{R}$. By definition:
$\Delta I(j,t;\mathcal{S}_{\nu})$ $=I(\mathcal{S}_{\nu})-\frac{n_{L}}{n_{\nu}}I(\mathcal{S}_{L})-\frac{n_{R}}{n_{\nu}}I(\mathcal{S}_{R})$
$=\left(\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}I(\mathcal{S}_{\nu,k})+\mathcal{E}(\nu)\right)$
$\quad-\frac{n_{L}}{n_{\nu}}\left(\sum_{k=1}^{K}\frac{n_{L,k}}{n_{L}}I(\mathcal{S}_{L,k})+\mathcal{E}(\nu_{L})\right)$
$\quad-\frac{n_{R}}{n_{\nu}}\left(\sum_{k=1}^{K}\frac{n_{R,k}}{n_{R}}I(\mathcal{S}_{R,k})+\mathcal{E}(\nu_{R})\right).$
Rearranging the terms and noting that $\frac{n_{L}}{n_{\nu}}\frac{n_{L,k}}{n_{L}}=\frac{n_{L,k}}{n_{\nu}}=\frac{n_{\nu,k}}{n_{\nu}}\frac{n_{L,k}}{n_{\nu,k}}$ (and similarly for the right child), we can group the local impurities for each client:
$\Delta I(j,t;\mathcal{S}_{\nu})$ $=\sum_{k=1}^{K}\frac{n_{\nu,k}}{n_{\nu}}\underbrace{\left(I(\mathcal{S}_{\nu,k})-\frac{n_{L,k}}{n_{\nu,k}}I(\mathcal{S}_{L,k})-\frac{n_{R,k}}{n_{\nu,k}}I(\mathcal{S}_{R,k})\right)}_{\Delta I(j,t;\mathcal{S}_{\nu,k})}$
$\quad+\underbrace{\mathcal{E}(\nu)-\left(\frac{n_{L}}{n_{\nu}}\mathcal{E}(\nu_{L})+\frac{n_{R}}{n_{\nu}}\mathcal{E}(\nu_{R})\right)}_{\Delta\mathcal{E}(j,t;\nu)}.$
This yields the decomposition in (5).
Finally, we show that for strictly concave impurities,
$\mathcal{E}(\nu)=0\iff\forall(k,k^{\prime}),\hat{\mathbb{E}}[Y\mid\mathcal{S}_{\nu},H=k]=\hat{\mathbb{E}}[Y\mid\mathcal{S}_{\nu},H=k^{\prime}].$
Let $\mu_{\nu,k}=\hat{\mathbb{E}}[Y\mid\mathcal{S}_{\nu},H=k]$ denote the local empirical mean outcome (regression) or class probability vector (classification).
For classification, $\mu_{\nu,k}$ is the vector of class probabilities (since $\hat{\mathbb{E}}[Y_{\text{one-hot}}]=\mathbf{p}$), and impurity is defined as $\Psi(\mu_{\nu,k})$ where $\Psi$ is strictly concave (e.g., Gini or Entropy). By Jensen’s inequality, the gap $\mathcal{E}(\nu)=\Psi\left(\sum\frac{n_{\nu,k}}{n_{\nu}}\mu_{\nu,k}\right)-\sum\frac{n_{\nu,k}}{n_{\nu}}\Psi(\mu_{\nu,k})$ is non-negative.
In both cases, due to strict concavity, $\mathcal{E}(\nu)=0\iff\mu_{\nu,1}=\dots=\mu_{\nu,K}$, i.e., the mean outcome or class distribution is identical across all clients.
C.4 Exact Computation of CART Criteria from Local Statistics
Proof of Theorem 3.5.
We show that standard CART impurities can be computed exactly from additive client-side sufficient statistics, without centralizing raw samples. Let $\mathcal{S}_{\nu,k}$ be the samples held by client $k$ at node $\nu$, with $n_{\nu,k}=|\mathcal{S}_{\nu,k}|$. Each client reports a statistic vector $\mathbf{s}_{\nu,k}$, and the server forms
$\mathbf{s}_{\nu}=\sum_{k=1}^{K}\mathbf{s}_{\nu,k}.$ (7)
The impurity at node $\nu$ is then obtained by a deterministic map $\Psi(\mathbf{s}_{\nu})$, which we now detail for the following standard CART criteria.
Regression: MSE.
Each client reports
$\mathbf{s}_{\nu,k}=\left(n_{\nu,k},\sum_{i\in\mathcal{S}_{\nu,k}}y_{i},\sum_{i\in\mathcal{S}_{\nu,k}}y_{i}^{2}\right).$ (8)
After aggregation, write $\mathbf{s}_{\nu}=(n_{\nu},S_{\nu},Q_{\nu})$. The MSE impurity is
$\Psi(\mathbf{s}_{\nu})=\frac{Q_{\nu}}{n_{\nu}}-\left(\frac{S_{\nu}}{n_{\nu}}\right)^{2}.$ (9)
Classification: Gini and entropy.
For $C$ classes, each client reports class counts
$\mathbf{s}_{\nu,k}=\left(n_{\nu,k},\sum_{i\in\mathcal{S}_{\nu,k}}\mathds{1}(y_{i}=1),\ldots,\sum_{i\in\mathcal{S}_{\nu,k}}\mathds{1}(y_{i}=C)\right).$ (10)
After aggregation, write $\mathbf{s}_{\nu}=(n_{\nu},N_{\nu,1},\ldots,N_{\nu,C})$. The Gini impurity is
$\Psi_{\mathrm{Gini}}(\mathbf{s}_{\nu})=1-\sum_{c=1}^{C}\left(\frac{N_{\nu,c}}{n_{\nu}}\right)^{2},$ (11)
and the Shannon entropy impurity is
$\Psi_{\mathrm{Ent}}(\mathbf{s}_{\nu})=-\sum_{c=1}^{C}\frac{N_{\nu,c}}{n_{\nu}}\log_{2}\left(\frac{N_{\nu,c}}{n_{\nu}}\right),$ (12)
with zero-count terms omitted.
Thus, for MSE, Gini, and entropy, the global CART impurity is a deterministic function of summed local statistics. Therefore, split gains can be evaluated exactly from federated aggregates alone. ∎
C.5 Top-L maximum gains aggregation recovers the true optimal feature
Proposition C.1 (Screening consistency under i.i.d. clients).
@@ C.5 Top-L maximum gains aggregation recovers the true optimal feature
Appendix D Impurity Decomposition under Covariate Shift
@@
Appendix E Exact Midpoint Enumeration
For completeness, we describe how centralized CART split candidates can be reproduced exactly in a federated setting. At a given node $\nu$, centralized CART evaluates midpoints between consecutive unique feature values. In principle, this can be achieved federatively by having each client transmit its node-level feature values, which the server merges into a globally sorted list.
In principle, this exact enumeration can be achieved in a federated architecture if every client $k\in\{1,\dots,K\}$ transmits its raw, local feature values for node $\nu$ to the server, which then merges and sorts them into a global list. Formally, for a given feature $j\in\{1,\dots,d\}$, let $\{v_{j,(r)}\}_{r=1}^{m_{\nu,j}}$ denote the sorted sequence of unique feature values pooled across all clients at node $\nu$, where $m_{\nu,j}\leq n_{\nu}$ represents the total number of unique values. The exact set of split candidates evaluated by the server is then:
$\mathcal{T}_{\nu}^{\mathrm{exact}}=\bigcup_{j=1}^{d}\left\{(j,t)\;\middle|\;t=\frac{v_{j,(r)}+v_{j,(r+1)}}{2},\quad 1\leq r<m_{\nu,j}\right\}.$ (14)
@@ Appendix E Exact Midpoint Enumeration
Appendix F Experimental Setting and Additional Simulations
@@
F.1 Covariate shift: when local objectives mis-rank global splits
@@ F.1.1 Disjoint supports
F.1.2 Overlapping supports: failure driven by the heterogeneity term
AvgImp can also be deceiving with overlapping supports, when the pooled impurity differs from the weighted average of local impurities due to the heterogeneity term $\mathcal{E}$ in Theorem 3.4.
To illustrate this, we simulate $K=2$ clients with features drawn as
@@ F.1.2 Overlapping supports: failure driven by the heterogeneity term
and a shared conditional model that is linear in the first feature:
@@
Under balanced mixture weights, the true global optimal split lies near the pooled population mean ($\mathbb{E}[X]=0$), a region well-populated by both clients. However, as the shift $\gamma$ increases, the heterogeneity term $\mathcal{E}(\nu)$ dominates. Consequently, maximizing the averaged local impurity gain incorrectly favors thresholds near the individual client means ($\pm\gamma$) rather than the global optimum.
Table 6 confirms this effect. While FedForest Quantiles consistently tracks the centralized baselines across all shift scales, AvgImp degrades sharply as $\gamma$ grows, proving that averaging local impurities is inherently vulnerable to covariate shift, regardless of support overlap.
Table 6: MSE comparison of methods with covariate shift.
@@
F.2 Covariate and outcome shifts
We consider the fully heterogeneous regime of Equation 3, where both covariates ($\alpha_{k}=.5$ and $\gamma_{k}=3$) and outcomes ($\delta=1.5$) vary across clients. Similar hyperparameters are used as in Section 5.